# Physically Feasible Semantic Segmentation

### v3

@@ 3.2 Computing Infeasible Inclusion Pairs
We apply Algorithm 1 to the training set of the dataset at hand for optimizing and evaluating the segmentation model, to compute the set $\mathcal{U}$ of class pairs corresponding to feasible, or valid, inclusions. Intuitively, this set contains class pairs for which a segment of the first class includes a segment of the second one at least in one case in the training set. The set of class pairs corresponding to infeasible inclusions, $\mathcal{C}^{\prime}$, is then simply the set-theoretic difference $\mathcal{C}^{\prime}=\mathcal{P}\setminus\mathcal{U}$, where $\mathcal{P}$ is the set of all possible non-identical class pairs.
In Algorithm 1, we first compute for each annotation $I$ in the training set two binary masks of the same dimension as $I$, in which pixels labeled as class $p$ and class $q$ are set as foreground, respectively. We then compute the connected components of class $q$, i.e. the class whose possibility to be included is checked, in mask_2. For each such connected component, we fill any holes inside it to remove any interior contour in the next step as it is irrelevant. After that, we compute the outer contour of the connected component at hand based on an 8-neighborhood and check whether it fully falls into pixels belonging to class $p$. If the latter is true and the length of the contour is non-negligible, we declare the case at hand as a valid inclusion and add the pair $(p,q)$ to the set $\mathcal{U}$.
@@ 3.3 Inclusion Loss
To compute our novel inclusion loss, we operate on the softmax maps $P_{\phi}(\mathbf{I})$ output by the segmentation network and index them channel-wise for each of the $|\mathcal{C}^{\prime}|$ infeasible inclusion pairs. Our core intuition is to identify in the pairs of softmax score maps that correspond to each infeasible inclusion type in $\mathcal{C}^{\prime}$ those regions across which class $c_{j}$—the includee—has a larger score than class $c_{i}$—the includer—and around which the opposite is true. These regions indicate “soft infeasible inclusions in the softmax predictions, which should be penalized proportionally to the area they occupy and to the variable “intensity of the inclusion, i.e. the magnitude of the difference between the softmax scores of classes $c_{i}$ and $c_{j}$ at each pixel of these regions. The computation of such a loss, which we term the inclusion loss, can be directly achieved in a differentiable formulation, amenable to standard backpropagation, via a grayscale morphological area opening of the aforementioned softmax difference maps. We implement this opening through an iterative algorithm, detailed in the following.
For each “infeasible pair $(c_{i},c_{j})\in\mathcal{C}^{\prime}$, we denote with $\text{P}_{\phi_{c_{i},c_{j}}}(\mathbf{I})\in\mathbb{R}^{(2\times H\times W)}$ the concatenation of $\text{P}_{\phi_{c_{i}}}(\mathbf{I})$ and $\text{P}_{\phi_{c_{j}}}(\mathbf{I})$. The difference between the two channels of the latter tensor is rectified with a ReLU as $\operatorname{ReLU}(\text{P}_{\phi_{c_{j}}}(\mathbf{I})-\text{P}_{\phi_{c_{i}}}(\mathbf{I}))\in\mathbb{R}^{(1\times H\times W)}$ (cf. Fig. 2). After that, all $|\mathcal{C}^{\prime}|$ pairwise differences are concatenated to obtain the tensor $\hat{P}_{\phi}(\mathbf{I})\in\mathbb{R}^{(|\mathcal{C}^{\prime}|\times H\times W)}$.
For each $(c_{i},c_{j})\in\mathcal{C}^{\prime}$, we initialize a running map $o^{(0)}_{(c_{i},c_{j})}\in\mathbb{R}^{(1\times{}H\times{}W)}$ so that its border pixels are set to 1 and the rest are equal to 0. The intuition here is that infeasibly included regions are always disconnected from the border of the image. Thus, positive values of the map $o^{(t)}_{(c_{i},c_{j})}$ that are gradually propagated from the border to the interior of the image via iterative 8-neighborhood-based dilation (a.k.a. max-pooling) will never reach the aforementioned regions if at each iteration they are multiplied with $\operatorname{ReLU}(\text{P}_{\phi_{c_{j}}}(\mathbf{I})-\text{P}_{\phi_{c_{i}}}(\mathbf{I}))$, as the latter is 0 everywhere around infeasibly included regions. The above, intuitively described iterative operation can be formally written as
@@
where $N(p)$ denotes the structuring element of the dilation operator or equivalently the max-pooling kernel, which we set to an 8-connected neighborhood, i.e. a 3x3 kernel. We iterate (2) $T$ times, where $T=\max\{2,\min\{H,W\}/2\}$ and $H$ and $W$ are the height and width of the input image. Namely, we iterate for the minimal number of times that is required for initial positive values at image borders in $o^{(0)}_{(c_{i},c_{j})}$ to be able to reach any pixel in the interior. The final map $o^{(T)}_{(c_{i},c_{j})}$ is by construction 00 at all pixels that belong to infeasibly included regions. This property implies that
$\frac{\hat{P}_{\phi,(c_{i},c_{j})}(\mathbf{I})\odot o^{(T)}_{(c_{i},c_{j})}}{\max\left\{o^{(T)}_{(c_{i},c_{j})},\epsilon\right\}}$ (3)
constitutes the grayscale area opening we are after for the infeasible inclusion class pair $(c_{i},c_{j})$ (cf. Fig. 3).
The relevant quantity for computing our inclusion loss is the difference between the original rectification output and its area opening in (3). This difference is given by
@@ 4.1 Datasets
Cityscapes.  Cityscapes [8] is a driving dataset, containing various road images of European cities. The dataset has 19 classes and includes 2975 training, 500 validation, and 1525 test images with fine annotations. Moreover, 20000 additional training images are coarsely annotated.
ADE20K.  We have used the dataset ADE20K [49]. It was earlier used in the ImageNet scene parsing challenge in 2016. There are 150 classes and diverse scenes with 1038 image-level labels. The dataset is divided into 20K/2K/3K images for training, validation, and test set.
ACDC.  The ACDC dataset [31] consists of driving scenes under adverse conditions such as fog, night, rain, and snow. It has 19 classes similar to Cityscapes. The dataset is divided into 1600 training, 406 validation, and 2000 test images.
@@ 4.2 Implementation Details
Input to PhyFea $\phi(I)$ is first interpolated to shape $\mathbb{R}^{(C\times 256\times 256)}$. The kernel size of the max-pooling used is $3\times 3$ with a stride of 1 and a padding of 1 to keep the dimensions the same and to check for any connected components not sharing the border pixels. For ADE20K, the number of infeasible inclusion pairs $|\mathcal{C^{\prime}}|$ is very large, so random sampling is applied offline in each iteration to reduce the memory footprint. PhyFea computes $l_{inclusion}$ on average for each iteration in 0.14 seconds. We report semantic segmentation performance using mean Intersection over Union (mIoU). The value of $\alpha$ (ref Eq. 1) is decided using a random search from the set [1e-5, 1e-10]. For all reported datasets, the value of $\alpha$ is taken as 1e-9. We have used 4 A100 GPUs each with memory 80GB.
Training setup.  For all the baselines, we have used a batch size of 4 for the Cityscapes [8] dataset and ACDC [31] datasets and 8 for the ADE20K [49] dataset. Optimizers, number of iterations, and learning rate parameters are kept similar to what is mentioned in the respective papers of the baselines for the respective datasets. For the ACDC [31] dataset, we applied the same setup as the corresponding baselines applied on the Cityscapes [8] dataset except for the crop size of the input.
We used mmsegmentation [7] to implement PhyFea on the baseline network SegFormer-B4 [41]. Random cropping of $1080\times 1080$ for the ACDC [31] dataset is used.
For OCRNet without SegFix [45], their official implementation is taken. For ACDC $769\times 769$ cropping is used.
For Mask2former [6], their official implementation is taken. For the ACDC dataset, we have taken a crop size of $540\times 960$.
Table 1: Comparison with state-of-the-art semantic segmentation methods on ADE20K, Cityscapes and ACDC. ’*’: multi-scale testing. Results are reported in mean IoU (%) on the Cityscapes test set, the ADE20K validation set, and the ACDC test set.
Method \ Dataset Backbone ADE20K Cityscapes ACDC
PSPNet [47] 44.4 78.5 -
DeepLabv2 [3] - 71.4 55.3
DeepLabv3+ [5] 44.1 80.9 70.0
HRNet [36] 43.1 80.9 75.0
@@
PhyFea w/ OCRNet (ours) HRNetV2-W48 46.8* 82.9* 67.6
@@
PhyFea w/ SegFormer (ours) MiT-B4 51.0 82.7 69.2
@@
PhyFea w/ Mask2Former (ours) Swin-L 56.7 84.0* 77.8
Table 2: Comparison of class-level IoU of OCRNet, SegFormer, and Mask2Former to their PhyFea-upgraded versions on the Cityscapes validation set. Training and evaluation are performed using the complete training and validation sets, respectively.
@@
Figure 4: Qualitative comparison on Cityscapes. From left to right: input image, ground-truth semantic labels, and predictions of Mask2Former [6] and PhyFea.
Table 3: Comparison of class-level IoU of OCRNet, SegFormer, and Mask2Former to their PhyFea-upgraded versions on the ACDC test set. Training and evaluation are performed using the complete training and test sets, respectively.
@@
Figure 5: Qualitative comparison on ADE20K. From left to right: input image, ground-truth semantic labels, and predictions of OCRNet [45] and PhyFea.
Figure 6: Training loss graph for Mask2Former and when re-trained with PhyFea on Cityscapes(top) and ACDC(bottom) datasets.
@@ 4.3 Comparison to The State of The Art
[…]
Table 2 is the class-wise comparison of IoU on Cityscapes val set for PhyFea and the corresponding baselines. We can observe that PhyFea achieved an overall improvement in IoU scores across the classes. PhyFea with OCRNet [45] shows improvement in 16 classes from the baseline, especially in the small classes like fence, rider, motorcycle, and bicycle which are very difficult concerning the semantic segmentation task. SegFormer [41] with PhyFea shows improvement in 14 classes from the baseline and despite performing poorly in the pole class, PhyFea has managed to acquire good improvement in other small and big classes. Mask2Former [6] shows improvement in 17 classes from the baseline. Mask2Former with PhyFea has the best improvement out of the other two baselines.
Table 3 is the class-wise comparison of IoU on the ACDC test set for PhyFea and the corresponding baselines. Here also we can observe that PhyFea achieved an overall improvement in IoU scores across the classes, especially over the SegFormer baseline. PhyFea achieved significant gains across all the baselines for the small classes like motorcycle, bicycle, and rider.
In Fig. 6 we can observe the training losses of baseline Mask2Former and when re-trained with PhyFea on the Cityscapes and ACDC datasets. At first, in both of the graphs we can see the loss with PhyFea is slightly higher than the baseline loss because, at first the $l_{inclusion}$ loss is very high up to 20k iterations roughly. The training loss tends to decrease after 20k iterations with PhyFea implementation slightly. Mask2Former implementation with PhyFea does not lose the ability to generalization on the test set. So we can confirm that Mask2Former with PhyFea is not over-fitting on the train set.
@@
Figure 7: Qualitative comparison on ACDC. From left to right: input image, ground-truth semantic labels, and predictions of baseline SegFormer network and PhyFea (ours).
Cityscapes results.  In Fig. 4, we note in the prediction of the basic SegFormer a segment of truck included by building, which is an infeasible inclusion. PhyFea, on the contrary, successfully opens (in morphological terminology) this truck segment. Moreover, in the SegFormer prediction, a segment of a motorcycle is completely included by a bicycle on the right of the image. PhyFea managed to completely open the motorcycle segment.
ADE20K results.  In Fig. 5, in the baseline prediction in the top row, a segment of vegetation is incorrectly included by another class, while PhyFea successfully opens the former segment. We can also observe that the overall condition of the prediction image after PhyFea re-training is not degraded from the baseline prediction. In the bottom row, a segment of the sea class is included by building class which is an infeasible inclusion. PhyFea managed to successfully open the included segment.
Table 4: Ablation Study performed on PhyFea with Mask2Former-SwinB on Cityscapes val set. Number of Iterations used for max-pooling Kernels $5\times 5$ and $7\times 7$ is 60k.
Kernel-size Stride Padding $\mathbf{T}$ mIoU
$3\times 3$ 1 1 128 85.3
$5\times 5$ 1 2 60 81.8
$7\times 7$ 1 3 3 79.8
ACDC results.  In Fig. 7 we can see in baseline prediction that some segments of the sidewalk class are included by the wall class. And PhyFea has successfully opened those segments. In the bottom row, we can observe that a wall segment is included by the building class which falls under the infeasible inclusion. PhyFea successfully opened the included segment in the final re-trained image.
4.4 Ablation Study
Tab 4 is the ablation study of PhyFea on the Cityscapes dataset using the Mask2Former as the baseline. We have shown the mIOU score on the val set. We can observe that by increasing the kernel size with a decrease in the number of iterations for max-pooling the mIOU score tends to decrease. We have kept the stride as 1 to compute for every possible position. We adjusted the padding appropriately to keep the dimensions the same for each iteration. The motivation behind the decrease in the mIOU score is that a bigger structuring element i.e. the max-pooling kernel, enables the propagation of the border pixels to the interior portion of the feature map which should otherwise be disconnected and as a result, the infeasible included segment will not be opened correctly. Thus, inclusion loss will be computed with a very high value penalizing non-infeasible included segments.

### v4

@@ 3.2 Computing Infeasible Inclusion Pairs
We apply Algorithm 1 to the training split of the dataset at hand for optimizing and evaluating the segmentation model, to compute the set $\mathcal{U}$ of class pairs corresponding to feasible, or valid, inclusions. Intuitively, $\mathcal{U}$ contains class pairs for which a segment of the first class includes a segment of the second one at least in one case in the training set. The set of class pairs corresponding to infeasible inclusions, $\mathcal{C}^{\prime}$, is then simply the set-theoretic difference $\mathcal{C}^{\prime}=\mathcal{P}\setminus\mathcal{U}$, where $\mathcal{P}$ is the set of all possible non-identical class pairs. In particular, any pair in $\mathcal{U}$ whose relative frequency of inclusion with respect to co-occurance of the class pairs is less than 2% is considered noise or an outlier, and is removed from $\mathcal{U}$.
In Algorithm 1, we first compute for each annotation $S$ in the training set two binary masks of the same dimension as $S$, in which pixels labeled as class $p$ and class $q$ are set as foreground, respectively. We then compute the connected components of class $q$, i.e. the class whose possibility to be included is checked, in mask_2. For each such connected component, we fill any holes inside it to remove any interior contour in the next step, as it is irrelevant. After that, we compute the outer contour of the connected component at hand based on an 8-neighborhood and check whether it fully falls into pixels belonging to class $p$. If the latter is true and the length of the contour is non-negligible, we declare the case at hand as a valid inclusion, and then if the relative frequency of the inclusion to the co-occurrence of the pair is non-negligible, we add the pair $(p,q)$ to the set $\mathcal{U}$.
@@ 3.3 Inclusion Loss
To compute our novel inclusion loss, we operate on the softmax maps $P_{\phi}(\mathbf{I})$ output by the segmentation network and index them channel-wise for each of the $|\mathcal{C}^{\prime}|$ infeasible inclusion pairs. Our core intuition is to identify in the pairs of softmax score maps that correspond to each infeasible inclusion case in $\mathcal{C}^{\prime}$ those regions across which class $c_{j}$—the includee—has a larger score than class $c_{i}$—the includer—and around which the opposite is true. These regions indicate “soft” infeasible inclusions in the softmax predictions, which should be penalized proportionally to the area they occupy and to the variable “intensity” of the inclusion, i.e. the magnitude of the difference between the softmax scores of classes $c_{i}$ and $c_{j}$ at each pixel of these regions. The computation of such a loss, which we term the inclusion loss, can be directly achieved in a differentiable formulation, amenable to standard backpropagation, via a grayscale morphological area opening of the aforementioned softmax difference maps. We implement this opening through an iterative algorithm, detailed in the following.
For each “infeasible pair” $(c_{i},c_{j})\in\mathcal{C}^{\prime}$, we denote with $\text{P}_{\phi_{c_{i},c_{j}}}(\mathbf{I})\in\mathbb{R}^{2\times H\times W}$ the concatenation of $\text{P}_{\phi_{c_{i}}}(\mathbf{I})$ and $\text{P}_{\phi_{c_{j}}}(\mathbf{I})$. The difference between the two channels of the latter tensor is rectified with a ReLU as $\operatorname{ReLU}(\text{P}_{\phi_{c_{j}}}(\mathbf{I})-\text{P}_{\phi_{c_{i}}}(\mathbf{I}))\in\mathbb{R}^{1\times H\times W}$ (cf. Fig. 2). After that, all $|\mathcal{C}^{\prime}|$ rectified pairwise differences are concatenated to obtain the tensor $\hat{P}_{\phi}(\mathbf{I})\in\mathbb{R}^{|\mathcal{C}^{\prime}|\times H\times W}$.
For each $(c_{i},c_{j})\in\mathcal{C}^{\prime}$, we initialize a running map $o^{(0)}_{(c_{i},c_{j})}\in\mathbb{R}^{1\times{}H\times{}W}$ so that its border pixels are set to 1 and the rest are equal to 0. The intuition here is that infeasibly included regions are always disconnected from the border of the image. Thus, positive values of the map $o^{(t)}_{(c_{i},c_{j})}$ that are gradually propagated from the border to the interior of the image via iterative 8-neighborhood-based dilation (a.k.a. max-pooling) will never reach the aforementioned regions if at each iteration they are multiplied with $\operatorname{ReLU}(\text{P}_{\phi_{c_{j}}}(\mathbf{I})-\text{P}_{\phi_{c_{i}}}(\mathbf{I}))$, as the latter is 0 everywhere around infeasibly included regions. The above, intuitively described iterative operation can be formally written as
@@
where $N(p)$ denotes the structuring element of the dilation operator or, equivalently, the max-pooling kernel, which we set to an 8-connected neighborhood, i.e. a 3x3 kernel. We iterate (2) $T$ times, where $T=\max\{2,\min\{H,W\}/2\}$ and $H$ and $W$ are the height and width of the input image $\mathbf{I}$. Namely, we iterate for the minimal number of times that is required for initial positive values at image borders in $o^{(0)}_{(c_{i},c_{j})}$ to be able to reach any pixel in the interior. The final map $o^{(T)}_{(c_{i},c_{j})}$ is by construction 00 at all pixels that belong to infeasibly included regions. This property implies that
$\frac{\hat{P}_{\phi,(c_{i},c_{j})}(\mathbf{I})\odot o^{(T)}_{(c_{i},c_{j})}}{\max\left\{o^{(T)}_{(c_{i},c_{j})},\epsilon\right\}}\vspace{-2mm}$ (3)
constitutes the grayscale area opening we are after for the infeasible inclusion class pair $(c_{i},c_{j})$ (cf. Fig. 3), where $\epsilon$ is a small positive constant.
The relevant quantity for computing our inclusion loss is the difference between the original rectification output and its area opening from (3). This difference is given by
@@ 4.1 Datasets
Cityscapes.  Cityscapes [11] is a driving dataset. The dataset has 19 classes and includes 2975 training, 500 validation, and 1525 test images with fine annotations and 20000 additional coarsely annotated training images.
ADE20K.  We have used the dataset ADE20K [52]. There are 150 classes and diverse scenes with 1038 image-level labels. The dataset is divided into 20K, 2K, and 3K images for training, validation, and testing, respectively.
ACDC.  The ACDC dataset [36] consists of driving scenes under adverse conditions such as fog, night, rain, and snow. It has the same 19 classes as Cityscapes. The dataset is divided into 1600 training, 406 validation, and 2000 test images.
@@ 4.2 Implementation Details
The softmax maps $P_{\phi}(\mathbf{I})$ input to InSeIn are first resized to $C\times 256\times 256$. The kernel size of the max-pooling used is $3\times 3$ with a stride of 1 and a padding of 1 to preserve spatial dimensions and to check for any connected components not sharing the border pixels. Random sampling on $|\mathcal{C^{\prime}}|$ is applied offline in each iteration to reduce the memory footprint. We report semantic segmentation performance using mean Intersection over Union (mIoU). The value of $\alpha$ is fixed to $10^{-9}$ for all examined datasets after a search in log-space in the interval $[10^{-10},10^{-5}]$ based on validation performance.
Training setup.  For all networks on which we implement InSeIn, we have used a batch size of 4 for Cityscapes and ACDC and 8 for ADE20K. The training protocol and hyperparameters remain the same as those for the base networks on each respective dataset. For ACDC, we applied the same setup as the corresponding baselines applied on the Cityscapes dataset. We used mmsegmentation [10] to implement InSeIn on the baseline network SegFormer-B4 [44].
Table 1: Comparison with state-of-the-art semantic segmentation methods on ADE20K, Cityscapes and ACDC. *: multi-scale testing. Results are reported on the Cityscapes test set, the ADE20K validation set, and the ACDC test set.
Method Backbone ADE20K Cityscapes ACDC
PSPNet [50] ResNet50 44.4 78.5 -
DeepLabv2 [6] ResNet-101 - 71.4 55.3
DeepLabv3+ [8] ResNet-101 44.1 80.9 70.0
HRNet [41] ResNet 43.1 80.9 75.0
@@
InSeIn w/ OCRNet (ours) HRNetV2-W48 46.8* 82.9* 67.6
@@
InSeIn w/ SegFormer (ours) MiT-B4 51.0 82.7 69.2
@@
InSeIn w/ Mask2Former (ours) Swin-L 56.7 84.0* 77.8
OneFormer [25] ConvNeXt-XL 57.4 83.8 78.0
InSeIn w/ OneFormer (ours) ConvNeXt-XL 58.0 84.3 78.9
SegMAN [15] SegMAN-B 52.6 83.6 77.5
InSeIn w/ SegMAN (ours) SegMAN-B 53.7 84.4 78.3
Table 2: Comparison to Beyond Pixels [22] on the Cityscapes test set using two baselines, SegFormer and Mask2Former. Beyond Pixels is also fully supervised, as InSeIn.
Method mIoU (%)
Beyond Pixels w/ SegFormer 82.5
InSeIn w/ SegFormer (ours) 82.7
Beyond Pixels w/ Mask2Former 83.7
InSeIn w/ Mask2Former (ours) 84.0
Table 3: Comparison of class-level IoU of OCRNet, SegFormer, Mask2Former, OneFormer, and SegMAN to their InSeIn-upgraded versions on the Cityscapes validation set. Training and evaluation are performed using the complete training and validation sets, respectively.
@@
OneFormer [25] 99.0 89.2 94.8 69.8 67.2 73.7 76.4 82.2 94.4 74.9 96.2 89.7 77.4 96.9 81.4 95.1 91.9 76.1 80.8 84.6
Ours 99.0 89.4 94.9 71.7 69.1 75.6 82.0 84.9 94.5 75.3 96.3 89.9 79.1 96.9 81.6 95.3 87.0 77.1 81.0 85.3
SegMAN [15] 98.9 88.4 94.3 65.2 65.9 72.8 79.5 82.9 94.3 74.3 96.1 88.6 75.9 96.6 79.3 93.8 91.5 74.8 79.7 83.8
Ours 98.8 88.3 94.3 67.9 67.8 73.5 80.6 83.9 94.3 74.4 96.0 89.2 75.8 96.8 83.6 94.1 91.2 74.0 80.0 84.5
Table 4: Comparison of class-level IoU of OCRNet, SegFormer, Mask2Former, OneFormer, and SegMAN to their InSeIn-upgraded versions on the ACDC test set. Training and evaluation are performed using the complete training and test sets, respectively.
@@
OneFormer [25] 94.5 80.9 91.9 66.8 57.7 64.3 74.9 75.0 79.2 68.7 87.7 78.0 62.1 92.9 84.9 91.5 94.0 68.7 68.4 78.0
Ours 98.7 86.9 93.3 58.8 60.3 65.7 73.0 78.3 93.5 72.8 95.6 85.9 71.2 95.9 73.3 82.3 69.5 67.2 75.9 78.8
SegMAN [15] 94.5 78.2 92.0 61.9 55.0 64.8 73.7 72.6 88.3 67.3 95.3 77.1 60.2 92.6 84.0 86.8 92.4 67.5 68.6 77.5
Ours 94.6 80.7 92.1 69.5 57.9 64.5 75.2 75.4 79.2 68.5 87.9 78.6 62.3 92.7 84.8 91.6 94.0 68.8 68.5 78.3
Figure 4: Qualitative comparison on Cityscapes. Best viewed on a screen with full zoom. From left to right: input image, ground-truth semantic labels, prediction of Mask2Former [9] network and InSeIn.
@@ 4.3 Comparison to The State of The Art
Tab. 1 compares InSeIn with a wide range of state-of-the-art methods for semantic segmentation on Cityscapes [11], ADE20K [52], and ACDC [36]. Overall, InSeIn achieves significant improvements in mIoU across all five network architectures on which we have implemented it and across all three datasets. On Cityscapes, we observe a consistent 0.50.50.50.50.80.80.80.8% improvement in mIoU with InSeIn on all architectures. On ADE20K, InSeIn achieves $1.2$% improvement over the baseline OCRNet [48] and $0.7$% over the baseline SegFormer-B4. On ACDC, InSeIn achieves $1.1$% improvement over the baseline OCRNet and $2.1$% over the baseline SegFormer-B4. Overall, improvements with InSeIn are consistent across all examined network architectures. This evidence shows that InSeIn can improve the segmentation performance of a general baseline network on which it is applied.
@@
Figure 5: Qualitative comparison on ADE20K. Best viewed on a screen with full zoom. From left to right: input image, ground-truth semantic labels, and predictions of OCRNet [48] and InSeIn.
Tab. 2 presents a comparison of InSeIn with a recent competing plug-and-play model, Beyond Pixels [22], on the Cityscapes dataset, implemented on two of our examined baseline networks, namely SegFormer and Mask2Former, for fully supervised segmentation. InSeIn consistently outperforms Beyond Pixels. Tab. 3 shows the class-wise comparison of IoU on the Cityscapes validation set for InSeIn and the corresponding baselines. We observe that InSeIn achieves a consistent improvement in IoU scores across most of the 19 classes. InSeIn with OCRNet [48], SegMAN [15], and SegFormer [44] show improvement in small challenging classes of the Cityscapes dataset. While we can observe an all-around and consistent improvement with Mask2Former [9] and OneFormer [25]. Tab. 4 shows the class-wise comparison of IoU on the ACDC test set for InSeIn and the corresponding baselines. We also observe that InSeIn achieves an overall improvement in IoU scores across the classes, especially over the SegFormer baseline. InSeIn achieves significant gains across all baselines for challenging small-instance classes such as motorcycle, bicycle, and rider.
Figure 6: Qualitative comparison on ACDC. Best viewed on a screen at full zoom. From left to right: input image, ground-truth semantic labels, and predictions of SegFormer [44] network and InSeIn.
Qualitative results on Cityscapes.  In Fig. 4, in the first row, baseline prediction, a building segment infeasibly includes a sky segment marked with green boxes, InSeIn rectifies it and also better segments the sidewalk marked with red boxes. The second row, baseline prediction, a bicycle infeasibly includes a motorcycle segment marked with green boxes on the right of the image. InSeIn rectifies this erroneous motorcycle segment.
Qualitative results on ADE20K.  In Fig. 5, in the first row, baseline prediction, the bridge segment includes a vegetation segment marked with red boxes, while InSeIn successfully rectifies such erroneous prediction. The second row, baseline prediction, building segment infeasibly includes a segment of the sea marked with green boxes. InSeIn rectifies it completely, and the red bounding boxes show additional parts of the prediction which are improved by InSeIn.
Qualitative results on ACDC.  In Fig. 6, in the first row, baseline prediction, wall segment infeasibly includes a sidewalk segment marked with green boxes. InSeIn successfully rectifies the infeasibility. The second row, baseline prediction, building segment infeasibly includes a wall segment marked with green boxes. InSeIn successfully predicts a wall label on that region. In both examples, red boxes highlight miscellaneous parts of the predictions where InSeIn improves upon baseline.
4.4 Evaluation of Prediction Feasibility
Here we demonstrate the efficacy of our model. The complete per-dataset counts of the feasible and infeasible inclusion pairs are reported in Tab. 5. Algorithm 1 is applied to the training set of every dataset, even when we train the networks on the union of the training and validation sets.
We define a class-level metric to measure how well predictions conform to our inclusion constraint, which we term mean infeasibility normalized frequency (mINF), as
$\centering\text{mINF}=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{(c_{i},c_{j})\in\mathcal{C}^{\prime}}\frac{f_{\text{inclusion}}(c_{i},c_{j})}{f_{\text{co-occur}}(c_{i},c_{j})}.\vspace{-2mm}\@add@centering$ (6)